OpenStudy (anonymous):

3^(2x+2)=8 (Exponential and Logarithmic Help)

OpenStudy (anonymous):

(2x+2)ln3=ln8 2x+2=ln8/ln3 2x=ln8/ln3-2 x=1/2*(ln8/ln3-2)

OpenStudy (anonymous):

Is the base 3?

OpenStudy (anonymous):

Hey, here's the solution to 9^(x+1) = 27 Let's express both sides to have the same base since 9 = 3^2 and 27 = 3^3 we can write the equation as 3^2(x+1) = 3^3 Having the same base we can equate the exponents, 2(x+1) = 3 you can now solve for x

OpenStudy (anonymous):

@Florcelda how did you get 27 rhs

OpenStudy (anonymous):

3 is not base just ln3

OpenStudy (anonymous):

Isn't this the first problem we tried to solve a while ago?

OpenStudy (anonymous):

3^(2x+2) = 8 (new problem) is it really a 3 raised to the power 2x + 2?

OpenStudy (anonymous):

can you verify the problem for me?